A length of copper of square cross section measuring 1.00mm by 1.00mm is stretched by a tension of 40N.What is the tensile stress in Pa?
1
Expert's answer
2020-08-28T09:33:10-0400
"P=\frac{F}{A}\\P=\frac{40}{0.001^2}=4.00\cdot10^7\ Pa"